### v1

# OCRBench v2: An Improved Benchmark for Evaluating Large Multimodal Models on Visual Text Localization and Reasoning

Abstract
Scoring the Optical Character Recognition (OCR) capabilities of Large Multimodal Models (LMMs) has witnessed growing interest recently. Existing benchmarks have highlighted the impressive performance of LMMs in text recognition; however, their abilities in certain challenging tasks, such as text localization, handwritten content extraction, and logical reasoning, remain underexplored. To bridge this gap, we introduce OCRBench v2, a large-scale bilingual text-centric benchmark with currently the most comprehensive set of tasks (4×4$\times$4 × more tasks than the previous multi-scene benchmark OCRBench), the widest coverage of scenarios ($31$ diverse scenarios including street scene, receipt, formula, diagram, and so on), and thorough evaluation metrics, with a total of $10,000$ human-verified question-answering pairs and a high proportion of difficult samples. After carefully benchmarking state-of-the-art LMMs on OCRBench v2, we find that $36$ out of $38$ LMMs score below $50$ ($100$ in total) and suffer from five-type limitations, including less frequently encountered text recognition, fine-grained perception, layout perception, complex element parsing, and logical reasoning. The benchmark and evaluation scripts are available at https://github.com/Yuliang-Liu/MultimodalOCR.
1 Introduction
Figure 1: Large multimodal models fail to deal with text-intensive tasks accurately. They are prone to errors in tasks such as text localization, handwritten content extraction, and mathematical reasoning, revealing limitations in tackling complex textual information within visual contexts.
The emergence of Large Language Models (LLMs) [1, 101, 8] has greatly improved the understanding and generation of structured text. However, in reality, much of the textual content is unstructured; it appears within images, videos, and other non-textual media in varied positions, orientations, and shapes. The need for processing such unstructured content leads to the study of Large Multimodal Models (LMMs) [5, 53, 139] that extend the text-only LLMs to additional modalities. By pretraining on multimodal data, LMMs acquire the zero-shot ability to interpret across diverse media such as recognizing and understanding complex visual scene text [59]. Such capability represents a significant advancement over standard Optical Character Recognition (OCR), because LMMs not only spot text but also interpret its semantic relevance to a scene.
Compared with classic OCR that typically relies on task-specific models to spot text, the increasing capability of LMMs to process and understand multimodal inputs has opened new potential to redefine the area of OCR. OCR has therefore become an important aspect of recent LMM evaluations. Some text-focused tasks have been included in standard benchmarks to assess the proficiency of LMMs in recognizing and interpreting textual content [26, 121]. Typically, text-based Visual Question Answering (VQA) datasets [93, 7, 107] are repurposed to evaluate OCR by framing generic VQA into questions that require accurate reading of embedded text. However, many of these text-centric datasets are initially created for classic OCR models, which is of limited diversity, depth, and suitability for evaluating LMMs. A common drawback is that, many questions lack sufficient complexity to assess the reasoning abilities of LMMs on scene text, and some can even be answered without visual input [10, 107].
Figure 2: Overview of the eight testable text-reading capabilities and associated tasks in OCRBench v2. Each color represents a distinct capability type.
More recently, several preliminary work [58, 46, 103, 50, 42] has attempted to explore the OCR capabilities of LMMs with customized benchmarks. For example, OCRBench [58] consolidates a range of common OCR tasks to evaluate LMM performance across traditional OCR functions. Other datasets, such as ComTQA [132] and ChartX [111], focus on structured text interpretation like table and chart understanding. While such effort represents a leap over standard OCR benchmarks, they remain limited in both data diversity and data quantity (see Table 1), which often leads to rapid performance saturation. For example, recent LMMs such as Qwen2-VL [105] have achieved 96.7% accuracy on the DocVQA dataset [71], nearly matching human performance at 98.1%, and 87.7% on OCRBench [58]. This raises an important question to the OCR community: Have OCR tasks truly been solved in the LMM era, or do existing benchmarks fail to capture the broader challenges present in diverse environments?
Benchmark #Scenario #Task #Image #Instruction
OCRbench [58] $\sim 14$ 5 0.9k 1k
Seed-bench-2-plus [46] $\sim 8$ 1 0.6k 2.3k
CONTEXTUAL [103] $\sim 11$ 1 0.5k 0.5k
Fox [50] 2 9 0.7k 2.2k
TableVQA-Bench [42] 1 3 0.9k 1.5k
MMTab-eval [134] 1 9 23k 49k
ComTQA [132] 1 4 1.6k 9k
ChartY [9] 1 1 6k 6k
ChartX [111] 1 7 6k 6k
MMC [51] 1 9 1.7k 2.9k
OCRBench v2 (Ours) 31 23 9.5k 10k
Table 1: Comparison between the proposed benchmark and existing text-centric datasets.
To answer the question above, we conducted preliminary tests with several state-of-the-art LMMs, including both open-source models such as Qwen2-VL-8B [105] and Monkey [48], and commercial models like GPT-4o [79]. These tests assessed performance on core OCR tasks, such as text localization, handwritten content extraction, and document-based logical reasoning. As illustrated in Figure 1, each model can fail on one of text-intensive tasks. These failures reveal a gap in detailed visual perception across different models, which constrains their effectiveness in tasks requiring accurate text localization, recognition, and contextual understanding within images. Such observations suggest that existing OCR benchmarks may be insufficient to assess the capabilities and limitations of LMMs for practical OCR applications due to limited task diversity, contextual complexity, and scale.
To bridge this gap, we propose OCRBench v2, a comprehensive benchmark designed to assess LMMs across a diverse set of OCR tasks. As shown in Figure 2, OCRBench v2 assesses eight core text-reading abilities, including text recognition, text referring, text spotting, relation extraction, element parsing, mathematical calculation, visual text understanding, and knowledge reasoning, organized into a total of $23$ concrete tasks. This benchmark provides $10,000$ high-quality, human-validated instruction-response pairs and also six types of evaluation metrics, which offers a rigorous framework for evaluating LMM performance in complex, practical OCR scenarios. In summary, the contributions of this work are three-fold:
  • OCRBench v2: an improved benchmark designed to evaluate the OCR capabilities within the context of LMMs. It assesses eight core OCR competencies and covers $23$ tasks across $31$ diverse scenarios ranging from scientific documents to natural scenes, which provides a thorough evaluation framework encapsulating fundamental and advanced text-centric challenges.
  • We systematically evaluate $38$ representative LMMs, ranging from commercial APIs to open-source models, which establishes broad baselines for OCR performance and enables a comparative understanding of model capabilities across varied OCR tasks.
  • We provide a detailed analysis to identify factors affecting the OCR capabilities of LMMs and highlight key insights for improvement. The analysis examines performance across various dimensions such as model generalization to diverse text types, model robustness, and the ability to tackle complex visual-textual relation.
2 Related Work
We review work related to OCR-enhanced LMMs and benchmarks for text-centric LMMs.
Figure 3: Sample visualizations for each task in OCRBench v2. The dataset provides a multilingual evaluation setting, comprising 23 sub-tasks grouped under 8 core OCR capabilities. Tasks marked with Refer to caption contain both English and Chinese instructions, while other tasks are either English-only Refer to caption or Chinese-only Refer to caption (Zoomed in for better clarity).
OCR-Enhanced LMMs.
Inspired by the success of LLMs, visual encoders are integrated into LLMs to create LMMs capable of processing both visual and textual information. Early LMMs exhibit strong zero-shot OCR capabilities, which motivates further exploration into text-centric LMMs. One pioneering approach, LLaVAR [131], enhances OCR abilities by training LLaVA [53] on text-centric datasets. mPLUG-DocOwl [118] constructs an instruction-tuning dataset with visual-text understanding to improve the model’s performance. However, these early methods are restricted to low-res inputs, limiting their ability to recognize dense and small text. To address this limitation, attention has shifted to increasing the input resolution of LMMs. DocPedia [25] processes high-res images in the frequency domain without increasing the input sequence length. UReader [119] crops an image into multiple sub-images, thereby achieving fine-grained perception of high-res text. On the contrary, LayoutLLM [67] adopts the LayoutLMv3 [33] as the vision encoder to process the high-res images. As the resolution of inputs increases, so do the challenges related to computational efficiency. To tackle this issue, TextMonkey [59] proposes a Token Resampler to compress redundant visual feature tokens, while mPLUG-DocOwl $1.5$ [31] introduces an H-Reducer to reduce the number of image tokens. As these models advance, they have achieved remarkable results on established benchmarks. Nevertheless, challenges remain unsolved in certain key areas such as text localization, entity extraction, and logical reasoning.
Benchmarks for Text-Centric LMMs.
To evaluate LMMs, developing a comprehensive benchmark is essential. Previous effort has focused on creating scenario-specific benchmarks to assess LMMs in particular contexts. For example, DocVQA [71] is designed to evaluate the document comprehension abilities of LMMs, while ChartQA [69] is tailored to chart interpretation skills. Similarly, Infographics VQA [73] is dedicated to assessing the understanding of infographic images. Additionally, TextVQA [93] aims to evaluate text comprehension in real-world scenes. To further investigate the robustness of the model, some methods expand the scope of evaluation scenarios. OCRBench [58] introduces a holistic evaluation framework that covers five core OCR tasks. CONTEXTUAL [103] is developed with context-sensitive instructions. SEED-Bench-2-Plus [46] encompasses a wide spectrum of text-rich images from various sources, including web content, maps, and charts. To provide a more thorough assessment, some benchmarks design multiple evaluation tasks within a specific scenario. TableVQA-Bench [42] first focuses on VQA tasks in the table domain. MMTab [134] and ComTQA [132] then extend the task scope, including table detection, structure recognition, and table querying. Moreover, ChartY [9], ChartX [111], and MMC [51] evaluate LMMs in chart understanding through tasks such as chart information extraction and reasoning. In this work, we focus on establishing a new benchmark called OCRBench v2, which contains more tasks than previous benchmarks and provides a systematic evaluation framework to reveal the limitations of LMMs in diverse text-rich environments.
3 OCRBench v2 Benchmark
In this section, we introduce OCRBench v2 from four aspects, including task, annotation, statistics, and evaluation.
3.1 Task Description
To provide a comprehensive evaluation framework for text-reading tasks, we categorize OCR capabilities into eight core areas, each encompassing specific sub-tasks that address various aspects of text comprehension and interpretation. Figure 3 demonstrates representative examples for each task, which showcases both visual inputs and corresponding instructions. Detailed descriptions of these core capabilities are as follows.
Text Recognition. This fundamental capability focuses on perceiving textual content. The related tasks include (fine-grained) text recognition and full-page OCR.
Text Referring. Determining the location of texts is necessary for real-world OCR applications. OCRBench v2 evaluates this ability with text grounding and VQA with position tasks, requiring LMMs to localize text positions accurately.
Text Spotting. Text spotting is a widely studied OCR task that requires models to output both the location and content of text. We consider it a distinct capability due to this unique output format.
Relation Extraction. Given that texts are often densely arranged in images, the ability to extract and map visual components is essential. This capability is assessed through key information extraction, key information mapping, and handwritten content extraction.
Element Parsing. Apart from text recognition, LMMs face the need of parsing complex elements for downstream applications. This ability is evaluated via table parsing, chart parsing, document parsing, and formula recognition.
Mathematical Calculation. Mathematical calculation abilities are essential for LMMs to address tasks that require numerical reasoning. Hence, text counting is introduced to assess the textual perception ability by counting the number of text instances. Additionally, we enhance the original math QA data by rendering textual questions into images, accompanied by geometric figures.
Visual Text Understanding. To tackle sophisticated tasks involving human interaction, LMMs need to comprehend semantic information of texts, a capability we term visual text understanding. This ability is evaluated by document classification and diagram QA. Additionally, we include basic VQA instructions where answers are located directly within the image, which refers to cognition VQA.
Knowledge Reasoning. Certain tasks require reasoning based on world knowledge and involve complex inference procedures, including science QA, APP agent interactions, ASCII art classification, and text translation. Further, we collect some complicated VQA instructions, where answers are not directly visible in images, termed reasoning VQA.
Due to space limitations, additional details for each subtask can be found in Appendix C.
Figure 4: The ratios of each type of tasks in OCRBench v2.
3.2 Annotation Curation
In this section, we introduce annotation curation in three parts, including dataset collection, instruction formulation, and manual verification.
Dataset Collection. To curate diverse data for OCRBench v2, we manually harvest and screen $81$ text-rich academic datasets. More details about the data sources can be found in Appendix B. To ensure diverse scenarios coverage, we also supplement them with additional private data. In all, our dataset comprises $31$ typical scenarios [^1] The $31$ scenarios include schematic diagrams, scientific papers, text image patches, filled tables, charts, receipts, question contexts, mathematical formulas, product labels, phone screenshots, indoor scenes, industry research reports, posters, street scenes, ASCII Art, shop signs, financial reports, chemical formulas, textbooks, magazines, emails, web screenshots, details page, verification code, resumes, illustrations, newspapers, road signs, menus, notify, and questionnaires., and the number of images in each scene is displayed in Table 9.
Instruction Formatting. To convert existing annotations into the LMM-compatible instruction format, we design specific prompts for each task. For complex tasks such as document parsing that require structured output, we include a format example to minimize the impact of instruction-following ability and focus the evaluation on OCR capabilities. Additionally, considering the distinct training strategies for localization tasks in LMMs, we standardize the coordinates by normalizing them with image sizes and by scaling to the range of $[0,1000]$. Such a standardization is explicitly specified in the related prompts. For unlabeled data, we manually annotate them according to task requirements.
Manual Verification. To ensure data quality, we manually review all instructions and correct approximately $1\%$ annotation errors.
Figure 5: OCR-related statistics of OCRBench v2.
Figure 6: Comparison of average entropy, type-token ratio, and variability index between OCRBench v2 and OCRBench.
Method Recognition Referring Spotting Extraction Parsing Calculation Understanding Reasoning Average
Open-source LMMs
Qwen-VL [5] 34.6 7.5 0 18.2 20.0 8.1 57.2 41.1 23.3
Qwen-VL-chat [5] 34.5 4.1 0 25.9 14.0 13.8 55.7 39.5 23.4
Qwen2-VL-8B [105] 72.1 47.9 17.5 82.5 25.5 25.4 78.4 61.5 51.4
InternVL2-8B [13] 49.9 23.1 0.5 65.2 24.8 26.7 73.5 52.9 39.6
InternVL2-26B [13] 63.4 26.1 0 76.8 37.8 32.3 79.4 58.9 46.8
InternVL2.5-8B [12] 59.0 25.0 1.4 77.5 35.1 29.4 75.3 57.2 45.0
InternVL2.5-26B [12] 65.6 26.1 1.6 86.9 36.2 37.4 78.3 62.9 49.4
TextMonkey [59] 39.1 0.7 0 19.0 12.2 19.0 61.1 40.2 23.9
LLaVA-Next-8B [52] 41.3 18.8 0 49.5 21.2 17.3 55.2 48.9 31.5
Monkey [48] 35.2 0 0 16.6 16.3 14.4 59.8 42.3 23.1
XComposer2-4KHD [21] 45.1 21.8 0.1 15.9 11.7 15.7 66.8 45.9 27.9
Molmo-7B [17] 52.4 21.3 0.1 45.5 7.6 28.5 65.3 55.0 34.5
EMU2-chat [95] 42.1 0.2 0 12.5 8.1 11.2 42.7 33.4 18.8
mPLUG-Owl3 [120] 41.6 14.0 0.6 24.4 10.9 11.1 52.2 46.0 25.1
CogVLM-chat [106] 50.9 0 0 0.2 8.4 15.0 58.1 41.7 21.8
Deepseek-VL-7B [62] 37.1 15.4 0 23.5 14.6 20.8 53.3 52.9 27.2
GLM-4V-9B [28] 61.8 22.6 0 71.7 31.6 22.6 72.1 58.4 42.6
MiniCPM-V-2.6 [117] 66.8 6.0 0.8 62.0 28.8 32.4 73.7 52.1 40.3
TextHarmony [133] 25.8 2.5 0 1.8 8.5 10.4 46.1 33.1 16.0
VILA1.5-8B [60] 35.3 15.5 0 21.1 12.7 17.3 46.3 40.3 23.6
LLaVAR [131] 37.3 0 0 1.0 9.9 12.3 34.6 27.0 15.3
DocOwl2 [32] 24.0 9.7 0 13.4 13.5 8.8 53.7 32.0 19.4
UReader [119] 22.4 0.1 0 0 9.2 7.9 41.0 29.1 13.7
Yi-VL-6B [122] 28.9 2.9 0 9.7 12.9 15.8 36.1 32.0 17.3
Janus-1.3B [110] 46.1 0 0 0.2 14.5 13.5 36.0 39.1 18.7
Cambrian-1-8B [100] 45.3 21.5 0 53.6 19.2 19.5 63.5 55.5 34.7
LLaVA-OV-7B [47] 46.0 20.8 0.1 58.3 25.3 23.3 64.4 53.0 36.4
Eagle-X5-7B [92] 34.7 17.8 0 21.7 20.6 21.5 61.0 42.6 27.5
Idefics3-8B [44] 23.8 13.2 0 63.2 23.8 23.0 65.8 44.9 32.2
Ovis1.6-3B [66] 59.2 14.3 0 65.0 32.1 29.0 69.8 56.8 40.8
Pixtral-12B [2] 48.9 21.6 0 66.3 35.5 29.8 66.9 53.7 40.3
Closed-source LMMs
GLM-4V-Plus [28] 60.3 25.2 0 74.7 37.6 26.4 61.4 57.2 42.9
GPT-4V [78] 69.7 26.9 0.3 75.6 36.7 42.9 71.5 57.9 47.7
GPT-4o [1] 61.2 26.7 0 77.5 36.3 43.4 71.1 55.5 46.5
GPT-4o-mini [77] 57.9 23.3 0.6 70.8 31.5 38.8 65.9 55.1 43.0
Gemini-Pro [99] 61.2 39.5 13.5 79.3 39.2 47.7 75.5 59.3 51.9
Claude3.5-sonnet [3] 62.2 28.4 1.3 56.6 37.8 40.8 73.5 60.9 45.2
Step-1V [94] 67.8 31.3 7.2 73.6 37.2 27.8 69.8 58.6 46.7
Table 2: Evaluation of existing LMMs on English subsets of OCRBench v2. “Recognition” refers to text recognition, “Referring” to text referring, “Spotting” to text spotting, “Extraction” to relation extraction, “Parsing” to element parsing, “Calculation” to mathematical calculation, “Understanding” to visual text understanding, and “Reasoning” to knowledge reasoning. Higher values indicate better performance. Best performance is in boldface, and the second best is underlined. The notations apply to all subsequent figures.
Method Recognition Extraction Parsing Understanding Reasoning Average
Open-source LMMs
Qwen-VL [5] 7.2 5.3 10.7 11.5 11.2 9.2
Qwen-VL-chat [5] 9.5 8.2 9.3 11.0 21.1 11.8
Qwen2-VL-8B [105] 51.3 51.4 21.6 52.5 37.5 42.9
InternVL2-8B [13] 20.6 45.2 23.2 54.4 38.1 36.3
InternVL2-26B [13] 21.9 46.0 34.8 50.9 34.8 37.7
InternVL2.5-8B [12] 52.8 52.8 28.6 56.4 40.5 46.2
InternVL2.5-26B [12] 32.4 56.1 32.6 56.3 43.6 44.2
TextMonkey [59] 23.5 14.8 8.4 19.9 12.2 15.8
LLaVA-Next-8B [52] 5.7 2.9 12.2 7.5 17.2 9.1
Monkey [48] 4.6 11.2 8.4 21.5 20.0 13.1
XComposer2-4KHD [21] 16.7 18.8 12.1 27.5 2.3 15.5
Molmo-7B [17] 7.1 15.0 9.2 9.0 23.7 12.8
EMU2-chat [95] 2.3 0.5 8.5 1.0 7.3 3.9
mPLUG-Owl3 [120] 6.6 17.9 9.7 6.0 26.1 13.3
CogVLM-chat [106] 5.5 10.0 9.8 1.5 2.5 5.9
Deepseek-VL-7B [62] 8.0 13.3 15.7 5.5 18.5 12.2
GLM-4V-9B [28] 24.4 60.6 20.4 52.8 25.2 36.6
MiniCPM-V-2.6 [117] 51.0 29.9 21.2 34.0 33.6 33.9
TextHarmony [133] 1.8 4.5 8.2 1.5 11.9 5.6
VILA1.5-8B [60] 5.4 8.8 8.5 3.0 15.5 8.2
LLaVAR [131] 2.3 1.7 8.9 0 2.5 3.1
DocOwl2 [32] 4.2 10.3 8.6 4.0 9.6 7.3
UReader [119] 6.8 2.7 8.4 2.5 7.2 5.5
Yi-VL-6B [122] 4.8 4.4 8.5 4.0 25.0 9.4
Janus-1.3B [110] 7.6 8.7 11.4 4.5 10.7 8.6
Cambrian-1-8B [100] 5.3 14.9 12.6 8.5 8.1 9.9
LLaVA-OV-7B [47] 14.8 15.7 13.7 16.0 28.7 17.8
Eagle-X5-7B [92] 7.5 12.0 11.6 5.0 19.2 11.1
Idefics3-8B [44] 7.0 15.5 15.9 9.0 18.1 13.1
Ovis1.6-3B [66] 11.5 23.7 22.8 28.8 18.9 21.1
Pixtral-12B [2] 13.4 10.9 21.0 7.0 20.7 14.6
Closed-source LMMs
GLM-4V-Plus [28] 34.5 60.6 23.9 49.8 28.2 39.4
GPT-4V [78] 49.9 52.2 34.6 40.8 22.9 40.1
GPT-4o [1] 21.6 53.0 29.8 38.5 18.2 32.2
GPT-4o-mini [77] 13.1 38.9 27.2 28.8 16.9 25.0
Gemini-Pro [99] 52.5 47.3 30.9 51.5 33.4 43.1
Claude3.5-sonnet [3] 21.0 56.2 35.2 55.0 30.5 39.6
Step-1V [94] 56.7 41.1 37.6 38.3 39.2 42.6
Table 3: Evaluation of existing LMMs on Chinese subsets of OCRBench v2.
3.3 Statistics of OCRBench v2
Here we present the statistics of OCRBench v2, including the proportion of eight task categories, the OCR-related statistics, and the measurement of prompt quality. More details on statistics can be found in Appendix D.
Capabilities Distribution. The proportion of eight capabilities in OCRBench v2 is shown in Figure 4. The proportions across different categories are relatively balanced.
OCR-Related Statistics. We further count the distribution of line-level OCR results of $7,400$ English and $2,600$ Chinese images in Figure 5(a) and Figure 5(b), respectively. The average number of line-level OCR results per category is shown in Figure 5(c). These statistics demonstrate that the text information is sufficiently rich in OCRBench v2.
Prompt Quality. We also compare the metrics of Average Entropy, Type-Token Ratio, and Average Variability Index of the questions between OCRBench v2 and OCRBench in Figure 6. Compared with OCRBench, OCRBench v2 shows greater Average Entropy, indicating higher unpredictability and diversity in its questions, which challenges the LMMs to tackle more diverse scenarios. It also exhibits a higher Type-Token Ratio, reflecting greater lexical variety and less redundancy in its question formulation. Additionally, the Average Variability Index of OCRBench v2 is significantly higher, suggesting a wider range of question types and structures, offering a more thorough performance evaluation of LMMs.
3.4 Evaluation Criteria
Since OCRBench v2 contains $23$ distinct tasks, a single evaluation metric is inadequate. Therefore, we adopt six types of evaluation metrics tailored to specific task categories. In the following, we present an overview of the evaluation metrics and their applicability to specific tasks. More details can be found in Appendix E.
Parsing Type. To evaluate the element parsing ability of LMMs, we assess their performance in transforming input images into structured formats, including HTML, Markdown, and JSON. TEDS [138] is employed to measure the structural similarity between outputs and the desired format.
Localization Type. For text referring, the IoU score is applied to quantify the distance between the predicted regions and the ground truth.
Extraction Type. To evaluate relation extraction, we employ the F1 score to assess key information extraction and mapping. Since this evaluation requires structural extraction of information from the output of LMMs, the format is provided in the given prompt.
Long Reading Type. To assess performance on tasks involving long text reading, BLEU [80], METEOR [6], F1 score, and edit distance are used to assess the similarity between predicted text and ground truth.
Counting Type. In text counting, LMMs are required to count the number of text instances. Thus, we use the L1 distance to measure the absolute difference between predicted and ground truth counts. The final score is then normalized to the range of $[0,1]$ based on the ground truth.
Basic VQA Type. To evaluate LMMs on general VQA cases, we employ task-specific metrics based on answer length and format. For questions where the original data provide options, we use exact string matching to compute accuracy. In other cases, we follow the approach of OCRBench to check whether the ground truth is contained in the prediction for short answers (fewer than $5$ words) and employ ANLS to measure prediction quality for longer answers ($5$ words or more).
4 Results and Discussions
Method Resolition Recognition Referring Spotting Extraction Parsing Calculation Understanding Reasoning Average
InternVL2-8B [13] 448 47.3 19.1 0.1 52.8 27.3 25.4 61.1 49.1 35.3
896 48.7 23.0 0.5 66.2 26.2 25.9 73.2 51.9 39.4
dynamic 49.9 23.1 0.5 65.2 24.8 26.7 73.5 52.9 39.6
Table 4: Evaluation of InternVL2-8B with different resolution setting on the English subsets of OCRBench v2.
Method Recognition Referring Spotting Extraction Parsing Calculation Understanding Reasoning Average
Qwen2VL-8B [105] 72.1 47.9 17.5 82.5 25.5 25.4 78.4 61.5 51.4
Qwen2VL-8B+OCR 69.8 50.4 20.1 79.1 29.4 28.0 77.7 60.0 51.8
Qwen2.5-8B+OCR 28.6 13.8 0 45.9 24.2 31.3 61.1 40.5 30.7
Table 5: Evaluation of Qwen2VL-8B and Qwen2.5-8B with pre-provided OCR information on English subsets of OCRBench v2.
Here we first benchmark state-of-the-art LMMs on OCRBench v2, presenting the quantitative analysis of their OCR capabilities. We then investigate several critical factors influencing OCR performance, including the resolution setting of visual encoder, pre-provided OCR information, and the direct integration between OCR information and language models. Further, we summarize key findings of current limitations and challenges for LMMs. All results are presented as percentages.
Figure 7: Results of eight representative LMMs across the eight ability dimensions defined in OCRBench v2.
4.1 Baselines
We evaluate a total of $31$ open-source LMMs including InternVL series [13, 12], Qwen-VL series [5, 105], GLM-4v-9B [28], Monkey [48], TextMonkey [59], LLaVA-Next-8B [52], XComposer2-4KHD [21], EMU2-chat [95], mPLUG-Owl3 [120], CogVLM-chat [106], Deepseek-VL-7B [62], Molmo-7B [17], MiniCPM-V-2.6 [117], TextHarmony [133], VILA1.5-8B [60], LLaVAR [131], DocOwl2 [32], UReader [119], Yi-VL-6B [122], Janus-1.3B [110], Cambrian-1-8B [100], LLaVA-OV-7B [47], Eagle-X5-7B [92], Idefics3-8B [44], Ovis1.6-3B [66], and Pixtral-12B [2] with the publicly available checkpoints. Further, some cutting-edge commercial models are also tested, including GPT4V [78], GPT4o [79], GPT4o-mini [77], Gemini-Pro [99], Claude3.5-sonnet [3], GLM-4V-Plus [28], and Step-1V [94].
4.2 Main Results
The main evaluation results are shown in Table 2, Table 3, and Figure 7. We observe that, while current LMMs perform relatively well on some basic capabilities such as text recognition and visual text understanding, most LMMs achieve low scores in other capabilities, such as text spotting and element parsing, mostly below $50$. In particular, some LMMs show significant limitations in text spotting capabilities, failing to precisely locate and recognize the texts. Additionally, LMMs demonstrate inadequate abilities in element parsing and mathematical calculation, which are crucial for complicated tasks like document analysis and mathematical reasoning. Besides, after comparing the performance of LMMs on visual text understanding and knowledge reasoning capabilities, we find that they perform poorly in knowledge reasoning, despite being provided with multiple-choice options. This suggests the deficiency of LMMs in logical reasoning.
4.3 Potential Factors Affecting OCR Capabilities
To provide additional insights, we also conduct experiments to explore the potential factors that may affect the performance of LMMs on OCRBench v2.
High-Res Visual Encoders. Since text often appears small in images, the resolution setting of the visual encoder could be a key factor affecting the text perception ability of LMMs [48]. Here we change the input resolution of the LMMs and observe the performance changes. In particular, InternVL2-8B is chosen, and the resolution setting includes $448$, $896$, and dynamic. Table 4 lists the results. Indeed, when the input resolution increases from $448$ to $896$, the performance increases by $4.1\%$.
Pre-provided OCR Information. To study the impact of OCR information, we use PaddleOCR [^2] https://github.com/PaddlePaddle/PaddleOCR to pre-extract OCR results and incorporate them with prompts. Table 5 shows the results. We observe that adding OCR information does not help much. This suggests that OCRBench v2 evaluates LMMs capabilities across multiple dimensions, rather than solely focusing on text recognition abilities.
Connection Between OCR and LLMs. We further explore a direct pipeline by first extracting OCR information and then by feeding it directly into Qwen2.5. Unlike LMMs, this pipeline separates OCR and language modeling into distinct stages. The results shown in Table 5 suggests that Qwen2VL-8B outperforms Qwen2.5 with OCR information, demonstrating LMMs’ remarkable ability to incorporate both textual and visual features efficiently.
4.4 Main Findings
Here we further highlight some interesting findings from the experimental results:
Finding 1. Limited recognition on less frequently encountered texts. Despite LMMs achieving high performance on text recognition, they can perform poorly when facing with less frequently encountered texts, such as dot matrix texts and mathematical formulas. This performance gap highlights the continuing challenges LMMs face in real-world text recognition.
Finding 2. Limited fine-grained spatial perception. Given that both closed- and open-source LMMs perform poorly in tasks such as text referring and text spotting in OCRBench v2, the current LMMs still lack the capability for accurate text localization.
Finding 3. Insufficient layout perception. While LMMs achieve good performance on basic text recognition, they struggle with complex layouts such as overlapping or rotated texts. For example, GPT-4o fails to detect all characters in overlapping handwritten text and misrecognizes numbers in 90° rotated images, revealing LMMs’ limitations in handling texts with complex layouts.
Finding 4. Weak analytics for complex elements. When facing complex visual elements like charts and formulas, LMMs struggle to parse them into structured data formats, hindering their effectiveness in downstream applications such as document digitalization.
Finding 5. Limited logical reasoning. While LMMs demonstrate basic multimodal reasoning abilities, they struggle with complex scenarios involving mathematical problems or complicated textual reasoning, highlighting the room for improvement in their reasoning capabilities.
5 Conclusion
In this work, we introduce OCRBench v2, an improved benchmark for assessing the OCR capabilities of LMMs. By studying $38$ representative LMMs across $23$ OCR tasks, we not only reveal some defects of the current LMMs but also explore factors affecting the OCR performance of LMMs. We hope OCRBench v2 could facilitate future research on improving the OCR capabilities of LMMs.
Limitations. One challenge we encountered is that LMMs sometimes produce responses that deviate from the given instructions, making it difficult to extract the desired answers. In future work, we plan to develop a more objective assessment framework to address this issue.
Acknowledgements
We thank Max Tyrrell for providing valuable receipt images, which have significantly contributed to the development of OCRBench v2.
Appendix
This appendix mainly contains the following contents:
  • Section A: Comparison experiments between LMMs and some text-centric expert models.
  • Section B: Data source of OCRBench v2.
  • Section C: Task definition of OCRBench v2.
  • Section D: More statistics of OCRBench v2.
  • Section E: Evaluation metrics of OCRBench v2.
  • Section F: Visualization samples of task examples.
  • Section G: Visualization samples of failure cases.
Appendix A Comparison with LMMs and Expert Models
Comparison with text recognizers. We compare LMMs with several representative scene text recognizers, including CRNN [87], ABINet [23], ASTER [89], MASTER [63], and SVTR [22], on the text recognition task. The weights of these models are loaded from mmocr [^3] https://github.com/open-mmlab/mmocr. The results are shown in Table 6, where we selected the top three LMMs with the best performance, including Qwen2VL-8B [105], GPT4V [78] and Step-1V [94]. The results demonstrate that LMMs exhibit remarkable text recognition capabilities, validating our motivation to evaluate LMMs on more challenging OCR-related tasks.
Method Accuracy
CRNN [87] 38.1
ABINet [23] 62.4
ASTER [89] 50.0
MASTER [63] 54.1
SVTR [22] 57.8
Qwen2VL-8B [105] 78.2
GPT4V [78] 79.1
Step-1V [94] 75.4
Table 6: Comparison between LMMs and text recognizers.
Comparison with text spotters. We also compare LMMs with ABCNet series [56, 57] and TESTR [130] on the text spotting task. The weights of the ABCNet series are loaded from mmocr and TESTR is loaded from the official checkpoint [^4] https://github.com/mlpc-ucsd/TESTR fine-tuned with TotalText [15]. The results are shown in Table 7. We observe that although LMMs demonstrate promising capabilities in text recognition, there remains significant potential for improvement in the text spotting task.
Comparison with GOT. We noticed a recent work, GOT [109], that can parse the textual elements within images. We conduct comparison experiments between GOT and some representative LMMs, and the results are shown in Table 8. We observe that LMM shows advantages in general text recognition, while GOT demonstrates better performance in document parsing task.
Method F1 score
ABCNet [56] 32.2
ABCNetV2 [57] 44.2
TESTR [130] 51.8
Qwen2VL-8B [105] 17.5
Gemini-Pro [99] 13.5
Step-1V [94] 7.2
Table 7: Comparison between LMMs and text spotters.
Method Rec FG-Rec Full-Rec Doc-Parse
GOT [109] 64.1 52.9 73.3 53.9
Qwen2VL-8B [105] 78.7 38.5 79.4 34.7
GPT4V [78] 79.5 20.0 80.4 39.2
Step-1V [94] 76.8 24.8 74.8 36.0
Table 8: Comparison between LMMs and GOT [109].
Appendix B Data Collection
Text Recognition. The data for text recognition task are sampled from ICDAR2013 [38], SVT [91], IIIT5K [75], ICDAR2015 [39], SCUT-CTW1500 [55], COCO-Text [102], CUTE80 [84], TotalText, SVTP [83], WordArt [112], NonSemanticText [58], IAM [68], ORAND-CAR-2014 [19], HOST [108] and WOST [108]. Meanwhile, CAPTCHA(Completely Automated Public Turing Test to Tell Humans Apart) images used in text recognition are sourced from a CAPTCHA dataset [^5] https://aistudio.baidu.com/datasetdetail/159309 and a number CAPTCHA dataset [^6] https://www.heywhale.com/mw/dataset/5e5e56b6b8dfce002d7ee42c/file. And dot matrix images in the text recognition task are obtained from the web page.
Fine-grained Text Recognition. In the fine-grained text recognition task, the images are sampled from the testset of Fox [50], Totaltext, COCO-Text, CTW1500 [127] and ICDAR2015. For the Fox dataset, we used the original annotations provided. For the other four datasets, we performed manual annotation.
Full-page OCR. The data sources for full-page OCR task include Fox, HierText [61], CTW [124], RCTW-17 [88], ReCTS [54], LSVT2019 [96], M6Doc [14], and CDLA [^7] https://github.com/buptlihang/CDLA.
Text Grounding. The images for the text Grounding task are sampled from testset of Totaltext, COCO-Text, CTW1500, and ICDAR2015. QA pairs and bounding boxes annotations are based on their official OCR annotations.
VQA with Position. The images for VQA with position task are sampled from testset of TextVQA [93] and RICO [18]. QA pairs and bounding boxes annotations are based on their official OCR annotations.
Text Spotting. The data sources for text spotting task include Totaltext, COCO-Text, CTW1500, and ICDAR2015.
Key Information Extraction. The data sources for key information extraction task include FUNSD [35], SROIE [34], POIE [43], M6Doc, XFUND [113], ICDAR2023-SVRD [123] and a private receipt photograph dataset.
Key Information Mapping. The data sources for the key information mapping task include FUNSD and POIE.
Handwritten Content Extraction. This task’s data is our private data, which contains real exam paper data with student information removed and manually annotated question-answer pairs.
Table Parsing. The images for table parsing task are selected from MMTab [134], WTW [85], TabRecSet [114] and flush table recognition competition [^8] https://github.com/10jqka-aicubes/table-recognition.
Chart Parsing. The data sources for the chart parsing task come from OneChart [9] and MMC [51].
Document Parsing. The data sources for document parsing task come from DoTA [49], DocVQA [72], M6Doc and CDLA.
Formula Recognition. The data sources for the formula Recognition task including HME100K [125], IM2LATEX-100K [104], M2E [115], MathWriting [27], MLHME-38K [^9] https://ai.100tal.com/icdar, CASIA-CSDB [20] and some private data.
Math QA. The data sources for the math QA task including MathMatics [86], MathVerse [129], MathVision [104] and MathVista [65].
Text Counting. The data for the text counting task are collected from IIIT5K, SVT, ICDAR2013, HierText, and TotalText.
Cognition VQA. The data sources for the cognition VQA task include EST-VQA [107], OCRVQA [76], ST-VQA [7], TEXTVQA, DIR300 [24], ChartQA [69], DVQA [37], PlotQA [74], InfoVQA [70], WTW, PubTabNet [136], WTQ [82], CORD [81], LLaVAR [131], WebSRC [11], DocVQA, M6Doc, XFUND, Publaynet [137], RVL-CDIP [29], ScreenQA [4], SlideVQA [97], a movie poster collection dataset [^10] https://www.kaggle.com/datasets/neha1703/movie-genre-from-its-poster, a website screenshot collection dataset [^11] https://huggingface.co/datasets/Zexanima/website_screenshots_image_dataset/tree/main and a private receipt photograph dataset.
Diagram QA. The data sources for the diagram QA task include AI2D [40] and TextBookQA [41].
Document Classification. The images for the document classification task are mainly collected from RVL-CDIP.
Reasoning VQA. The reasoning VQA task shares some common data sources with the cognition VQA task. Additionally, portions of the reasoning VQA dataset are drawn from MMSI [45] and CMMMU [128].
Science QA. The images and annotations of the science QA task are collected from ScienceQA [64] and MMMU-Pro [126]
APP Agent. The data source of the APP agent task is RICO.
ASCII Art Classification. The data sources for the ASCII art classification task is ASCIIEval [36].
Text Translation. The datasets collected for text translation task includes memes [^12] https://www.kaggle.com/datasets/dvishal485/meme-challenge?resource=download, MSRA-TD500 [116], MTWI2018 [30], M6Doc, ICDAR2023-SVRD, EST-VQA, RCTW17 [90], DAST1500 [98], XFUND, ArT2019 [16], ChartQA, CDLA, ICDAR2015, SlideVQA, Fintabnet [135], ScienceQA, InfoVQA, COMICS-Dialogue [^13] https://huggingface.co/datasets/lmms-lab/M4-Instruct-Data and ExpressExpense SRD [^14] https://expressexpense.com/blog/free-receipt-images-ocr-machine-learning-dataset/.
Appendix C Task Definition
In this section, we introduce the definition of each task, and the visualizations for each task can be found in Section F.
Text Recognition. Text recognition refers to the fundamental OCR ability on text image patches, which asks LMMs to read the text content. To comprehensively evaluate LMMs’ text recognition ability across diverse scenarios, our collection incorporates various text types, including regular text, irregular text, artistic text, handwriting text, digit string text, non-semantic text, occluded text, doc matrix text, and CAPTCHA text.
Fine-grained Text Recognition. This task requires LLMs to read and comprehend textual content within the given region. It evaluates LLMs’ fine-grained perception capabilities in understanding text in natural scenes and documents.
Full-page OCR. Full-page OCR [50] task requires LMMs to extract and recognize all text content from the given images. Converting text into digital format facilitates subsequent processing and analysis of text images.
Text grounding. In this task, users would provide a text string and require LMMs to locate its specific location, evaluating LMMs’ fine-grained perception capabilities.
VQA with Position. For VQA with position task, LMMs need to not only respond to the question but also provide the exact position coordinates that directly correspond to the answer. We ask LMMs to output both information in JSON format for convenient evaluation, and the coordinates are required to be normalized with image sizes and scaled to the range of $[0,1000]$.
Text Spotting. Text spotting task needs LMMs to output the localization and content of all appeared text simultaneously. Due to the interference of background elements and the large number of text instances, this task demands high fine-grained perception capabilities from the model. Besides, the coordinates are required to be normalized with image sizes and scaled to the range of $[0,1000]$.
Scene Number Scene Number Scene Number
Schematic diagram 1238 Scientific paper 799 Word 728
Table(filled) 705 Chart 620 Receipts 609
Questions 581 Mathematical formula 475 Product labels 434
Phone screenshot 431 Indoor scenes 395 Industry research reports 343
Poster 264 Street scene 224 ASCII Art 199
Shop sign 189 Financial reports 153 Chemical formula 149
Textbook 148 Magazine 146 Email 111
Web screenshot 99 Details page 95 Verification code 87
Resumes 67 Illustration 61 Newspaper 52
Road signs 43 Menus 31 Notify 30
Questionnaire 29
Table 9: The number of images included in each scene category.
Figure 8: The quantity distribution of English tasks.
Figure 9: The quantity distribution of Chinese tasks.
Figure 10: The OCR lines distribution of English tasks.
Figure 11: The OCR lines distribution of Chinese tasks.
Key Information Extraction. The key information extraction task is to extract the necessary information from densely arranged text. In this task, we provide some desired entities as keys and demand LMMs to output the corresponding values to form the output JSON string.
Key Information Mapping. In this task, we provide a set of entity keys and their corresponding values in the prompt. The LMMs are then asked to match and pair these keys with their respective values into groups.
Handwritten Content Extraction. To investigate the information extraction capabilities of LMMs in educational scenarios, we collect some Chinese examination papers, containing both printed question text and handwritten student responses. There are four types of questions in these examination papers, including single-choice, multiple-choice, true or false, and brief response questions. The prompts require LMMs to extract the handwritten content for specific questions.
Table Parsing. Table parsing task requires LMMs to parse the given table into structured text, including Markdown and HTML format.
Chart Parsing. Apart from tables, charts can also be converted to structured information. In this task, LLMs are required to transform visual charts into JSON format.
Document Parsing. In the document parsing task, both text and the complex elements, including chart, table, and formula, are required to parse.
Formula Recognition. This task asks LMMs to recognize the given formula in the latex format. The collection includes mathematical and chemical formulas.
Math QA. Math QA task evaluates the LMMs’ mathematical calculation ability. In particular, we render the mathematical problem description and related figures into images and ask LMMs to answer the questions within the images.
Text Counting. Text counting task is built to evaluate the quantity property perceiving ability of LMMs, including the character frequency in words and the word counting in the given image.
Cognition VQA. In OCRBench v2, we split text-centric VQA instructions into cognition VQA and Reasoning VQA based on whether the answers can be directly found in the images. Cognition VQA task refers to the instructions where answers are explicitly present in the given image. This task evaluates the fundamental text-centric question-answering ability based on visual content.
Diagram QA. In the diagram QA task, LMMs need to respond to the question about the given diagrams, reflecting LMMs’ ability to understand the relationship between the visual elements.
Document Classification. Document classification task asks LMMs to classify the category of the given document image. The included categories are letters, forms, emails, handwritten documents, advertisements, scientific reports, scientific publications, specifications, file folders, news articles, budgets, invoices, presentations, questionnaires, resumes, and memos.
Reasoning VQA. In reasoning VQA tasks, the answers often do not directly appear in the image. This forces LMMs to perform logical reasoning to respond to questions based on visual information.
Science QA. In the Science QA task, LMMs are required to respond to the scientific problem. We use PaddleOCR [^15] https://github.com/PaddlePaddle/PaddleOCR to extract text from the collected images and filter out those with fewer than four OCR results. Additionally, when extra subject-related knowledge is provided by the source, we incorporate them by rendering them into the images.
APP Agent. For the APP agent task, LMMs need to understand the relationship between textual content, icons, and world knowledge to respond to the question from the user, simulating the real-world application scene.
ASCII Art Classification. We incorporate a recent image classification task that uses images composed purely of ASCII characters [36]. This task is included in OCRBench v2 to evaluate LMMs’ ability to assess LMMs’ pattern recognition and visual abstraction abilities.
Text Translation. In the text translation task, LMMs need to execute translation between Chinese and English texts, evaluating LMMs’ semantic understanding abilities.
Appendix D More Statistics of OCRBench v2
D.1 Scene Coverage
Our dataset can be divided into 31 classic scenes according to the scene of the image. The specific scenes and the corresponding number of pictures are shown in Table 9.
D.2 Statistics of each task
The distribution of 23 tasks in OCRBench v2 is displayed in Figure 8 and Figure 9. Additionally, we calculate and present the average number of OCR text lines per task in Figure 10 and Figure 11. As illustrated in these figures, the task distribution is well-balanced, with each task containing adequate textual information for analysis.
Appendix E Evaluation Metrics
Parsing Type. We use Tree-Edit-Distance-based Similarity (TEDS) [138] to evaluate parsing tasks, which require LMMs to transform the images to structured formats. Tree Edit Distance (TED) refers to the minimum number of edits to transform one tree into another. TEDS is based on TED to calculate the similarity of two trees. Assuming $T_{1}$ and $T_{2}$ are two different trees, $TED(T_{1},T_{2})$ refer to their TED, the TEDS is defined as:
$TEDS(T_{1},T_{2})=1-\frac{TED(T_{1},T_{2})}{\max(|T_{1}|,|T_{2}|)},$ (1)
where $|T_{1}|$ and $|T_{2}|$ is the number of nodes of trees, $TED(T_{1},T_{2})$ can be calculated by dynamic programming algorithm. If $T_{1}$ and $T_{2}$ are identical, then their TEDS equals to $1$. As the structural difference between two trees increases, their TED value becomes larger, resulting in the TEDS approaching 00.
Localization Type. In the text referring and spotting tasks, LMMs are required to provide regression bounding boxes of target objects. IoU score is adopted to measure the distance between the predicted regions and the ground truth.
$IoU(B_{1},B_{2})=\frac{Intersect(B_{1},B_{2})}{Union(B_{1},B_{2})},$ (2)
where $Intersect(B_{1},B_{2})$ refers to the overlap area of bounding box $B_{1}$ and $B_{2}$, while $Union(B_{1},B_{2})$ refers to their union area.
Extraction Type. The F1 score is used to evaluate LMMs’ relation extraction capability. Given the predicted and ground truth Key-Value pairs, the F1 score is formulated as follows:
$Precision=\frac{N_{3}}{N_{2}},$ (3)
$Recall=\frac{N_{3}}{N_{1}},$ (4)
$Fmean=\frac{2*Precision*Recall}{Precision+Recall},$ (5)
where $N_{1}$, $N_{2}$, and $N_{3}$ denote the number of ground-truth Key-Value pairs, predicted Key-Value pairs, and correctly matched Key-Value pairs, respectively.
Long Reading Type. To evaluate LMMs’ ability to recognize text across entire paragraphs or pages, BLEU [80], METEOR [6], F1 score and normalized edit distance are employed. And the final score is the average value of these metrics.
BLEU evaluates prediction quality by comparing n-gram match rates between prediction and ground truth sequences. For each n-gram type, precision is calculated as the ratio of matching n-grams to total prediction n-grams. The final BLEU score is the geometric mean of these precision values multiplied by a penalty $BP$, which is defined as:
$BLEU=BP*exp(\sum^{N}_{n=1}{w_{n}\log{p_{n}}}),$ (6)
$BP=\begin{cases}1&L_{p}\geq L_{g}\\ e^{(1-\frac{L_{p}}{L_{g}})}&L_{p}<L_{g}\end{cases},$ (7)
where $p_{n}$ represents the precision of n-grams, $L_{p}$ represents the length of prediction sequence, $L_{g}$ represents the length of ground truth sequence, $w_{n}$ is weight factor, usually evenly distributed ($w_{n}=\frac{1}{N}$). Typically, $N$ is set to 4.
METEOR employs a semantic-aware matching strategy with four levels. 1) Exact Match: words in prediction that are identical to ground truth. 2) Stem match: matching words that have the same word stem. 3) Synonym Match: matching words based on synonymous relationships. 4) Paraphrase Match: Matching similar phrases at the phrase level. These matches are combined to calculate precision and recall, from which a weighted harmonic mean F1 score is derived as:
$P_{meteor}=\frac{N_{match}}{N_{pred}},$ (8)
$R_{meteor}=\frac{N_{match}}{N_{gt}},$ (9)
$F_{meteor}=\frac{10*P_{meteor}*R_{meteor}}{P_{meteor}+9*R_{meteor}},$ (10)
where $N_{match}$, $N_{pred}$, and $N_{gt}$ represent the number of matched items, words in prediction, and words in ground truth, respectively. The final METEOR score is obtained by multiplying the $F_{meteor}$ by the penalty adjustment factor. The calculation is formulated as follows:
$METEOR=F_{meteor}*(1-BP_{meteor}),$ (11)
$BP_{meteor}=0.5*\frac{N_{chunk}}{N_{match}},$ (12)
where $N_{chunk}$ refers to the number of contiguous matching phrases. More chunks indicate greater word order differences, resulting in a heavier penalty.
The calculation method of the F1 score in long reading metrics follows the same approach as discussed in extraction metrics, as shown in Equations 3, 4, 5.
Normalized Edit Distance (NED) measures string similarity by computing the minimum number of operations needed to transform one string into another. And then NED is normalized by the length of the longer string. The calculation is formulated as follows:
$NED(S_{1},S_{2})=\frac{ED(S_{1},S_{2})}{\text{max}(len(S_{1}),len(S_{2}))}$ (13)
where $ED(S_{1},S_{2})$ represents the edit distance between the prediction string $S_{1}$ and the ground truth $S_{2}$. The $NED$ value of 0 indicates identical strings, while 1 indicates completely different strings.
Counting Type. In OCRBench v2, character frequency counting and word counting tasks are included. For character frequency, we use exact match evaluation since the answers are typically single-digit integers. For word counting, we evaluate using the L1 distance between predicted and ground truth counts, normalized to $[0,1]$ based on the ground truth. This can be formulated as follows:
$score=\begin{cases}0&C_{pred}\leq 0\\ 1-\frac{|C_{pred}-C_{gt}|}{C_{gt}})&0<C_{pred}<2*C_{gt}\\ 0&C_{pred}\geq 2*C_{gt}\\ \end{cases},$ (14)
where $C_{pred}$ and $C_{gt}$ denote the predicted count and ground truth count, respectively.
Basic VQA Type. The remaining tasks in OCRBench v2 are basic VQA types, and we employ different evaluation metrics based on question types. For multiple-choice questions, we use exact match between predictions and answer options. In other cases, we check whether the ground truth is contained in the prediction for answers shorter than $5$ words, and use ANLS for longer answers.
Appendix F Samples for Each Task
As show in Figure 12 to Figure 20 , there are 23 OCR tasks included in OCRBench v2. Among them, Figure 12 to Figure 18 present examples of English tasks, including text recognition, diagram QA, text counting, formula recognition, math QA, VQA with position, ASCII art classification, reasoning VQA, text translation, APP agent, table parsing, cognition VQA, document classification, science QA, chart parsing, key information extraction, full-page OCR, text spotting, fine-grained text recognition, text grounding, key information mapping, and document parsing. These figures show corresponding images and QA pairs for each of the 22 tasks. Figure 19 to Figure 20 provide examples of Chinese tasks, including key information extraction, text translation, formula recognition, reasoning VQA, cognition VQA, handwritten content extraction, document parsing, full-page OCR, and table parsing, along with their associated images and QA pairs.
Figure 12: Samples for each tasks.
Figure 13: Samples for each tasks.
Figure 14: Samples for each tasks.
Figure 15: Samples for each tasks.
Figure 16: Samples for each tasks.
Figure 17: Samples for each tasks.
Figure 18: Samples for each tasks.
Figure 19: Samples for each tasks.
Figure 20: Samples for each tasks.
Appendix G Samples for LMMs’ Limitations
Figure 21 to Figure 23 provide examples corresponding to the findings discussed in Section 4.4 of the main text, which show error results of GPT-4o [1], Monkey [48], and Qwen2VL-8B on various tasks in OCRBench v2. These examples highlight the current limitations of LLMs on OCR tasks. For instance, LLMs exhibit poor recognition of less frequently encountered texts, struggle to accurately locate text in tasks involving text and coordinates, and demonstrate insufficient perception of text in complex layouts such as rotated texts. Additionally, their logical reasoning abilities are limited when addressing mathematical problems, and their analysis of complex elements in charts remains weak. These are the capabilities of LLMs in OCR tasks that require further improvement.
Figure 21: Samples for LMM’S Limitations. The portion of LLM’s response marked in red is incorrect content, and the content in the red dashed box is missing information.
Figure 22: Samples for LMM’S Limitations. The portion of LLM’s response marked in red is incorrect content, and the content in the red dashed box is missing information.
Figure 23: Samples for LMM’S Limitations. The portion of LLM’s response marked in red is incorrect content, and the content in the red dashed box is missing information.